A body of mass 0.5 kg travels in a straight linr with velocity v=ax3/2 where a = 5m-1/2s-1. what is the work done by the net force during its displacement from x=0 to x=2m?

Grade:11

1 Answers

Susmita
5 years ago
v=ax3/2
differentiate wrt x.
dv/dt=(3a/2)x1/2(dx/dt)=(3a/2)x1/2v=(3a2/2)x2
force F=mdv/d
Work done dW=Fdx
Integrate from 0 to 2.
W=(a2m/2)8=4a2m=4*(52)*0.5=50J